a) If I wanted to set up a controlled breeding experiment to determine the inheritance pattern of the Vermillion eye color in Drosophila and determine if that trait was inherited under the Mendelian Ratio of 3:1, what steps would I need to conduct for the experiment and how would I monitor that experiment and analyze the data? That is, how would I set up that experiment (briefly) and what data am I collecting to analyze and with what statistical test?

b) Describe how to use the Chi-Square statistical analysis for determining if my experimental data fit the mendelian ratio I was testing with my experimental hypothesis.

c) Describe 4 reasons Drosophila are especially useful as model organisms in conducting controlled breeding experiments?
Answer #1

Solution a

Steps for the experiment

  1. Selecte pure breeding Drosophila for Vermillion eye color
  2. Cross the pure breeding Drosophila for Vermillion eye color
  3. Select the F1 progeny and do selfing
  4. Count the number of F2 offsprings. - Take total and seperate count of Homozygous dominant, Heterozygous dominant and Homozygous recessive
  5. You may repeat the above steps - Replications
  6. We know the expected phenotype ratio is 3:1 (suppose if we got a total of 200 offspings, 150 would show dominant trait and 50 would show recessive trait)

Solution b.

For doing the Chsquare analysis, follow the formula

- )-*

where, O is the observed number of phenotypes and E is the expected number of phenotypes

Suppose in your experiment you got 150 pogenies in the F2 and out of the 200, 80 are Homozygous dominant, 60 are Heterozygous dominant (Total 140 dominants) and 60 are Homozygous recessive, these are the observed observed numbers. By using Chisquare we have to analyse whether there numbers folloe the real (expected) ratio of 3:1 (dominant:recessive)

For example if we follow the above number, the Chisquare analysis will be like this,

Phenotype

Observed

Expected

O-E

(O-E)2

(O-E)2/E

Dominant

140

200*(3/4)= 150

140-150 = -10

100

100/150 = 0.67

Recessive

60

200*(1/4)= 50

60-50 = 10

100

100/50 = 2

Total

200

Ʃ = 2.67

The Chisquare value we obtained is Ʃ = 2.67.

Now we have to compare this Chisquare value (calculated chisquare value) with the tabled Chisquare value (It is a contant value) at 1 degrees of freedom and 0.05 probability level . The tabled value at 1 degrees of freedom and 0.05 probability level is 3.841

calculated chisquare value we obtained is 2.67 and tabled value is 3.841.

Here at 1 degrees of freedom, and 0.05 probability level the calculated value (2.67) of χ² is lesser than tabled value (3.841) Therefore the Null hypothesis is accepted.

Our Null Hypothesis is that 'there is no significant difference in the observed ratio and expected ratio (3:1)' and our re chisquare analysis shows that there is no there is no significant difference in the observed ratio and expected ratio and the experimental results follow Mendelian ratios.

If the calculated value of χ² is greater than tabled value (3.841) we will reject the Null hypothesis and rejection means the results do not follow mendelian ratios.

Solution c

Reasons to use Drosophila as model organism

  1. Simple and well understood genome
  2. Short life cycle
  3. Quick reproduction rate
  4. Easy to grow under in-vitro conditions
  5. Drosophila genome is 60% homologous to human genome